Find the cutoff wavelength for the continuous X-rays coming from an X-ray tube operating at 30 kV.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
V = 30 Kv λ = hc/E = hc/ev = 1242 ev – nm/e * 30 * 10^3 = 414 * 10^-4 nm = 41.4 Pm.